A bead sliding on a wire has a speed of 210 cm/s at A. If friction forces are negligible, what will

be the speed at pt. B and C?



1
Expert's answer
2021-04-09T06:49:59-0400

To be given in question

Bead velocity in point A "v_{A}=2.1meter\/sec"

To be asked in question

Point B and pointC Speed find out

"v_{B}=?;v_{C}=?"

We know that energy of conservation


at point B speed

"\frac{1}{2}m{v}_f^2+\frac{1}{2}m{v}_i^2+mg(h_{f}-h_{i}) =0"

At point B hight zero

"h_{f}=0"

"\frac{1}{2}m{v}_f^2-\frac{1}{2}m{v}_i^2+mg(0-h_{i}) =0"


"v^2_{f}-v^2_{i} =2gh_{i}"

"v_{f}^2=2gh_{i}+v^2_{i}"

"v_{f}=\sqrt {2gh_{i}+v^2_{i}}"


Put value "v_{i}=v_{A}=2.1meter\/sec"

"v_{f}=\sqrt{2gh+4.41} meter/sec"

Point B velocity "v_{B}=v_{f}=[v_{f}]_{B}"

"[v_{f}]_{B}=\sqrt{2gh+4.42} meter/sec"

Pont c velocity

"\frac {1}{2}mv^2_{f}-\frac{1}{2}mv^2_{i} +mgh_{f}-mgh_{i}=0"

"v^2_{f}-v^2_{i} =2g(h_{i}-h_{f})"

"v^2_{f}=v^2_{i} +2g(h_{i}-h_{f})"

"v_{f}=\sqrt {v^2_{i} +2g(h_{i}-h_{f})}"

"v_{f}=v_{c}=[v_{f}]_{c}"

"[v_{f}]_{c}=\sqrt{4.41+2g(h_{i}-h_{f}) }meter/sec"